OpenStudy (anonymous):

I need help! I'm taking my online exam! A circle has a radius of 3 cm. if an angle has a measurement of 2 raidians what is the length of the arc that corasponds to the angle?

OpenStudy (anonymous):

arc= 1/2 r * teta = 1/2 * 3* 2 = 3

OpenStudy (anonymous):

Thank you so much ! ☺️

OpenStudy (anonymous):

i think 1/2 r^2 * tetaa= please wait?

OpenStudy (anonymous):

1/2 * 9 * 2 = 9